OpenStudy (anonymous):

Find the slope and y-intercept of each line. y = x-3

OpenStudy (anonymous):

slope: 1, y-intercept: -3?

OpenStudy (anonymous):

yes, correct slope intercept form is y = mx + b where m is the slope and b is the y-intercept

OpenStudy (anonymous):

what about this one? 4. Transform this equation to slope-intercept form: 6x + 4y = 8

OpenStudy (anonymous):

4y = 8 - 6x transposing the x-term to other side no we divide all terms by 4 and get 3 y = 2 - ---- x 2 or y = - 3/2x + 2 so slope is -3/2 and y-intercept is 2

OpenStudy (anonymous):

thanks again @Harkirat

OpenStudy (anonymous):

u r welcome...☺